OpenStudy (anonymous):

Which of the following is equivalent to log9w? log9 + logw log9 - logw w(log9) 9(logw)

OpenStudy (anonymous):

@welshfella

rishavraj (rishavraj):

\[\log a + \log b = \log (ab)\] \[\log a - \log b = \log (\frac{ a }{ b })\]

rishavraj (rishavraj):

\[\log (a^x) = x \log a\]